OpenStudy (shamil98):

ln(e^(3)x^(4)) ln e^3 + ln x^4 = 3 + ln x^4 ?

OpenStudy (phi):

and ln(x^4) = 4 ln(x) so 3 + 4 ln(x)

OpenStudy (shamil98):

Oh, okay. thank you again.

OpenStudy (phi):

the idea is, for example, ln(x^3) = ln(x*x*x) = ln(x) + ln(x) + ln(x) =3 ln(x)

OpenStudy (shamil98):

So it's like saying. \[\large \ln~a^b = \ln~ba\]

OpenStudy (shamil98):

\[\large b~ \ln a\] correction*